# Collisional whistler instability and electron temperature staircase in inhomogeneous plasma

### v1

@@ 4 Linear stability condition
Less well-understood is the second stabilization term due to ‘resistivity curvature’ [^2] In Davies et al. (2021), it is discussed how $\partial_{z}\eta$ can also act as an advection velocity of the magnetic field, analogous to $v_{N}$, so one might also expect that $\partial_{z}^{2}\eta$ would enter as the pile-up of this advection velocity. But if this effect were present, the sign of the resistivity-curvature drive would be opposite to that seen in (35b) and there would be a corresponding Doppler shift in (35a). Hence, this cannot be the physical origin of the resistivity-curvature drive.. As shown in Fig. 7, the diffusion of a perturbation is faster when $\partial_{z}^{2}\eta>0$ than homogeneous theory would predict (thus increasing the stability of the system against the perturbation), and the diffusion is slower when $\partial_{z}^{2}\eta<0$ (decreasing the stability of the system). Per Fig. 2, one generally has $\partial_{z}^{2}\eta>0$ in the vicinity of a hotspot, with $\partial_{z}^{2}\eta$ becoming increasingly larger as the magnetization level decreases.
@@
$\frac{\beta_{0}\Delta_{\beta}^{2}(\mathcal{M})}{\alpha_{\perp}(\mathcal{M})}\geq\frac{15\alpha_{\perp}(\mathcal{M})}{\beta_{0}\mathcal{M}^{2}}-\frac{15\alpha_{\perp}^{\prime}(\mathcal{M})}{\beta_{0}\mathcal{M}}+\frac{25\alpha^{\prime\prime}_{\perp}(\mathcal{M})}{\beta_{0}}-10\beta_{\wedge}^{\prime}(\mathcal{M}).$ (49)
Note that (49) is actually independent of the temperature gradient. Hence, when the right-hand side is sufficiently positive there will be no unstable temperature gradients. It is clear that this will happen for weakly magnetized plasmas (small $\mathcal{M}$) or for low-$\beta_{0}$ plasmas due to the divergent denominator [^3] The plot of (49) as a function of $\mathcal{M}$ and $\beta_{0}$ is nearly identical to the plot of $F_{T,1}$ in Fig. 9 for reasons that will be discussed in Sec. 5, with the unstable region colored in red and the stable region colored in blue.; this weakly-magnetized regime will be discussed further in Secs. 6 and 7.
@@ 5 Dynamical relevance of collisional whistler instability
Note that here and in what follows we use to denote $\partial_{\mathcal{M}}$ for univariate functions of $\mathcal{M}$.
@@
with dynamical relevance requiring $\gamma_{\text{whist}}\tau_{\kappa}>1$. Note that when $S=-5/2$ the diffusion time becomes infinite ($T\propto z^{2/7}$ yields a spatially constant heat flux), so the whistlers will be dynamically relevant anywhere there is a positive growth rate.
@@
First of all, there is no visible difference between the results for an isobaric versus a constant density profile. This is because the density-drive terms in (50), $F_{n,1}$ and $F_{n,2}$, are generally smaller than the principal temperature-drive term $F_{T,1}$ (Fig. 9). Secondly, not all of the parameter space is susceptible to whistlers even when the initial profile is diffusion-free ($S=-2.5$). This is because the instability actually disappears for sufficiently low $\mathcal{M}$ and all whistler waves are instead strongly damped. This is in stark contrast to the prediction made with the short-wavelength asymptotic growth rate obtained in Bell et al. (2020):
@@ 6 Global marginally stable temperature profiles
We shall first discuss the general case before considering two special cases in detail. The first special case has the plasma density constrained by pressure balance, as occurs in astrophysical and recent experimental contexts (Markevitch & Vikhlinin, 2007; Meinecke et al., 2022). The second special case will be the simpler situation in which the density is constant, corresponding to the analysis performed in Bell et al. (2020).
@@ 6.1 General theory
Let us seek monotonic profiles such that $\partial_{z}\mathcal{M}\neq 0$ everywhere, $\mathcal{M}$ having been defined in (17) but with $\epsilon=0$. Then one can formally parameterize the inverse function $z(\mathcal{M})$ so that all functions can be considered functions of $\mathcal{M}$. Suppose further that
@@
where $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are arbitrary values at which boundary conditions can be applied. One notes that (70) manifestly respects the affine symmetry of the original equation. For (70) to be physically relevant, though, it must be the case that $\mathcal{M}(z)\geq 0$; a sufficient condition to ensure positivity is that $g\sim A/\mathcal{M}$ with $A>1$ as $\mathcal{M}\to 0$, as shown in Appendix E. Realistic friction coefficients do indeed have this property, as we shall see in Sec. 6.3 and Sec. 6.4.
@@ 6.2 Magnetization staircases as a general class of solutions
Let us now consider a high-$\beta_{0}$ plasma. Although not obvious from (70), in this limit, $\mathcal{M}(z)$, and thus $T(z)$, naturally forms a staircase structure. To see this more easily, note that $G_{1}\sim O(\delta^{-1})$ and $G_{2}\sim O(1)$ with respect to the small parameter $\delta\sim 1/\beta_{0}$ (more simply, $F_{T,1}$ is significantly larger than any other coefficient in Fig. 9 when $\beta_{0}$ is large). Hence, (67) has the general abstract form
@@
where $G$ is a nominally $O(1)$ function. It is well established (Bender & Orszag, 1978) that the solutions to such equations can exhibit boundary layers when $\delta\to 0$; for (71) such boundary layers will occur where $G(y)=0$.
@@
Figure 11: Solutions of (71) when $G(y)$ takes the form shown in each respective inset. All solutions have $|\delta|=0.01$. The ‘arbitrary units’ (a.u.) designation on the $x$-axis emphasizes that, due to affine symmetry, there is formally no scale to the $x$-dependence of the solutions. All solutions satisfy $y(0)=0$, with the other boundary condition $y^{\prime}(0)$, which simply controls the horizontal scale of the solution, adjusted for each case to fit the pertinent behaviour on the same axis. The functional forms for the plots shown in (a) and (b) can be derived analytically, as shown in Appendix F.
The basis for this conjecture is demonstrated in Fig. 11, which shows solutions of (71) for a polynomial $G(y)$. The derivation of (72) and the analytical solution for certain special cases of $G(y)$ are presented in Appendix F. Let us now demonstrate the role that these staircase solutions play in determining the globally stable temperature profiles for isobaric and constant-density plasmas.
@@ 6.3 Isobaric density profile with Chapman–Enskog friction
Using known asymptotics of the Lorentz friction coefficients (Appendix B), it is straightforward to show that $\mathcal{M}g\to 8/5$ as $\mathcal{M}\to 0$. Hence, the temperature profile is guaranteed to be positive everywhere (Appendix E). To obtain a simple analytical approximation for the solution to (70), it is reasonable to take
@@
Note, importantly, that a geometrical-optics description of this parameter regime is not valid because (43) predicts that $k_{z,\text{max}}L\sim\mathcal{M}_{*}\ll 1$ when $\mathcal{M}_{*}^{2}\beta_{0}\sim 1$ and $\mathcal{M}_{*}\ll 1$. The continuation of the root line to small $\beta_{0}$ can be computed using the $\mathcal{M}\to\infty$ limit of the friction coefficients (Appendix B) to give
@@
As shown in Appendix F, the solution to (70) can be computed analytically for $g(\mathcal{M})$ given by (78):
@@ 7 Back-reaction on the background temperature profile
The appropriate lowest-order expressions for $\eta$, $v_{N}$, and $\mathcal{D}_{A}$ are given in (27) and (35b). It is important to note that, when combined with (36), (90b) manifestly conserves the total energy of the electron MHD system of equations (the appropriate expressions for energy conservation are presented in Appendix G).
@@ 7.1 Frictional cooling
$-\frac{\mathcal{M}\beta_{0}}{2\alpha_{\perp}}\left(\sqrt{\beta_{\wedge}^{2}+\Delta_{\beta}^{2}}+\beta_{\wedge}\right)<L_{T}L_{I}^{-1}<\frac{\mathcal{M}\beta_{0}}{2\alpha_{\perp}}\left(\sqrt{\beta_{\wedge}^{2}+\Delta_{\beta}^{2}}-\beta_{\wedge}\right).$ (96)
@@ 7.2 Reduced heat flux
$-\frac{q_{z}}{q_{0}}=\kappa_{\parallel}+\frac{\epsilon^{2}}{2}\left(\frac{\Delta_{\beta}^{2}}{2\alpha_{\perp}}+\frac{\beta_{\wedge}}{\mathcal{M}\beta_{0}}\frac{L_{T}}{L_{I}}-\Delta_{\kappa}\right)\mathcal{I},\quad q_{0}=\frac{nTv_{t}^{2}\tau_{ei}}{L_{T}}.$ (97)
Hence, we see that the net effect of the instability on the heat flux results from the competition of three terms. The first two $O(\epsilon^{2})$ terms are associated with the Ettingshausen effect [^6] For detailed discussions of the Ettingshausen effect, see, e.g., Chittenden & Haines (1993) and Kolmes et al. (2021a)., which is the additional heat flux [beyond the standard enthalpy flux (Epperlein & Haines, 1986)] carried by faster moving, less collisional electrons whose directional symmetry is broken with a mean flow. The first of these terms is always positive and therefore always enhances the heat flux; in contrast, the second term can reduce the heat flux when $L_{I}$ and $L_{T}$ are oppositely oriented and can even overcome the first term if $L_{T}L_{I}^{-1}$ is sufficiently negative (meaning that $\mathcal{I}$ is sufficiently sharply peaked):
@@
where the final approximation is for $\mathcal{M}\ll 1$. This heat-flux-reduction mechanism can be readily achieved in the high-$\beta_{0}$, $-\mathcal{M}$ regime in which the temperature staircases discussed in Sec. 6 also form, since in this limit the right-hand side of (98) goes to zero as $-25/\beta_{0}$ [see (79)]. The third $O(\epsilon^{2})$ term in (97), which is always negative, is the reduction of the effective conductivity due to the transverse magnetic field perturbations generated by the instability causing the temperature gradient and the total magnetic field to become misaligned.
@@ 7.3 Marginally stable heat flux
Finally, let us suppose that the temperature profile is in a globally marginally stable state such that $\left\langle\mathcal{D}_{A}\right\rangle=0$ (Sec. 6). Dynamically, one expects an arbitrary temperature profile to be driven towards such a state on the instability timescale, which can be faster than the conduction timescale (see Fig. 10). This is because, as shown in Fig. 9, the instability growth rate $\mathcal{D}_{A}$ is an increasing function of temperature (and conversely, the damping rate is a decreasing function of temperature). Energy conservation (90b) then implies that higher temperatures are increasingly cooled by the instability (thereby decreasing $\mathcal{D}_{A}$) while lower temperatures are increasingly heated (thereby decreasing $-\mathcal{D}_{A}$) to create temperature plateaus separated by transition regions where the temperature profile has remained unchanged because $\mathcal{D}_{A}\approx 0$ initially. The result is a profile that has $\mathcal{D}_{A}=0$ everywhere.
@@
$-\frac{\mathcal{Q}_{z}}{\kappa_{\parallel}q_{0}}=1-\epsilon^{2}\Gamma\mathcal{I},$ (99)
@@
Here we have imposed pressure balance for simplicity; the general expression is obtained by replacing $5/2$ with $L_{T}\partial_{z}\mathcal{M}/\mathcal{M}$ in the second term. As in (97), we see that the intensity gradient is capable of reducing the total heat flux. In this case, the two mechanisms available are the Ettingshausen mechanism discussed in Sec. 7.2 and the frictional cooling discussed in Sec. 7.1 (now coupled to the heat flux by imposing global marginal stability). Both are controlled by the lengthscale ratio $L_{T}L_{I}^{-1}$. As shown in Fig. 15, making $L_{T}L_{I}^{-1}$ negative causes the heat flux to be reduced ($|\mathcal{Q}_{z}|<\kappa_{\parallel}q_{0}$) over a large region of parameter space777 One can have reduced heat flux even with $L_{T}L_{I}^{-1}>0$ provided that $\beta_{0}$ and $\mathcal{M}$ are both sufficiently large ($\beta_{0}\mathcal{M}\gtrsim L_{T}L_{I}^{-1}$) such that $\Delta_{\kappa}$ dominates (100)..
@@
Indeed, the flux-reduction factor $\Gamma$ can be made arbitrarily large by making $L_{T}L_{I}^{-1}$ increasingly negative [^8] Of course, having $L_{T}L_{I}^{-1}$ too large will cause the perturbation approach underlying (91) to break down (even an infinitesimal intensity will grow quickly in space to be relatively large), but as an academic exercise, we can nevertheless study what the constraints would be on $\mathcal{I}(z)$ to greatly reduce the heat flux in the idealized scenario that quasilinear theory holds for all $L_{T}L_{I}^{-1}$. Future investigations can be conducted to see how nonlinear physics modifies this constraint.. To make this more quantitative, let us consider the weakly magnetized ($-\mathcal{M}$) limit and approximate all transport coefficients by their lowest-order asymptotic limits (Appendix B). Then one has
@@
or, equivalently, using (101) [^9] The inequality (104) represents a stricter requirement on $L_{I}$ compared to (98) because (104) requires second-order effects to become comparable to the lowest-order heat flux, whereas (98) results from comparing two second-order terms.,
$-\frac{L_{T}}{L_{I}}\geq\frac{\beta_{0}}{0.423}\left(\frac{\kappa_{\parallel}+\kappa_{\perp}}{2\kappa_{\parallel}}\right)\gtrsim 2\beta_{0},$ (104)
@@
Thus, the collisional whistler instability is capable of reducing the electron heat flux in principle, somewhat similar to the collisionless whistler instability (Levinson & Eichler, 1992; Pistinner & Eichler, 1998; Komarov et al., 2018; Roberg-Clark et al., 2018). The persistence time $\tau$ of the temperature profile is then nominally lengthened by the same factor $\tau\sim\tau_{\kappa}/|1-\epsilon^{2}\Gamma\mathcal{I}|$ until it is ultimately set by the persistence time of the intensity profile itself, which in turn is set by advection and refraction (i.e., evolving the $z$ and the $k_{z}$ dependence of $\mathcal{W}$, respectively). The advection-limited persistence time is expected to be comparable to $\tau$ (and thus not a limiting factor) since the total Poynting flux that accounts for the intensity advection [the term in square brackets in (157)] and the modified heat flux $\mathcal{Q}_{z}$ used to estimate $\tau$ only differ by subdominant terms. The refraction timescale, however, is difficult to estimate given that it inherently involves breaking the quasi-eikonal ansatz (92), requiring one to reconsider the full phase-space dynamics of $\mathcal{W}$. This is beyond the scope of the present work.

### v2

@@ 5 Linear stability condition
Less well-understood is the second stabilization term due to ‘resistivity curvature’. As shown in Fig. 7, the diffusion of a perturbation is faster when $\partial_{z}^{2}\eta>0$ than homogeneous theory would predict (thus increasing the stability of the system against the perturbation), and the diffusion is slower when $\partial_{z}^{2}\eta<0$ (decreasing the stability of the system). Per Fig. 2, one generally has $\partial_{z}^{2}\eta>0$ in the vicinity of a hotspot, with $\partial_{z}^{2}\eta$ becoming increasingly larger as the magnetization level decreases.
@@
$\frac{\beta_{0}\Delta_{\beta}^{2}(\mathcal{M})}{\alpha_{\perp}(\mathcal{M})}\geq\frac{15\alpha_{\perp}(\mathcal{M})}{\beta_{0}\mathcal{M}^{2}}-\frac{15\alpha_{\perp}^{\prime}(\mathcal{M})}{\beta_{0}\mathcal{M}}+\frac{25\alpha^{\prime\prime}_{\perp}(\mathcal{M})}{\beta_{0}}-10\beta_{\wedge}^{\prime}(\mathcal{M}),$ (45)
where as a reminder, $\mathcal{M}$ is defined in (17) with $\epsilon=0$. Note that (45) is actually independent of the temperature gradient. Hence, when the right-hand side is sufficiently positive there will be no unstable temperature gradients. It is clear that this will happen for weakly magnetized plasmas (small $\mathcal{M}$) or for low-$\beta_{0}$ plasmas due to the divergent denominator [^2] The plot of (45) as a function of $\mathcal{M}$ and $\beta_{0}$ is nearly identical to the plot of $F_{T,1}$ in Fig. 9 for reasons that will be discussed in Sec. 6, with the unstable region colored in red and the stable region colored in blue.; this weakly-magnetized regime will be discussed further in Secs. 7 and 8.
@@ 6 Dynamical relevance of collisional whistler instability
Note that here and in what follows, we use to denote $\partial_{\mathcal{M}}$ for univariate functions of $\mathcal{M}$. Also note that all the length scales in (47) are signed quantities.
@@
with dynamical relevance requiring $\gamma_{\text{whist}}\tau_{\kappa}>1$. Note that when S=5/2S=-5/2 the diffusion time becomes infinite ($T\propto z^{2/7}$ yields a spatially constant heat flux), so the whistlers will be dynamically relevant anywhere there is a positive growth rate.
@@
First of all, there is no visible difference between the results for an isobaric versus a constant density profile. This is because the density-drive terms in (46), $F_{n,1}$ and $F_{n,2}$, are generally smaller than the principal temperature-drive term $F_{T,1}$ (Fig. 9). Secondly, not all of the parameter space is susceptible to whistlers even when the initial profile is diffusion-free (S=2.5S=-2.5). This is because the instability actually disappears for sufficiently low $\mathcal{M}$ and all whistler waves are instead strongly damped. This is in stark contrast to the prediction made with the short-wavelength asymptotic growth rate obtained in Bell et al. (2020):
@@ 7 Global marginally stable temperature profiles
We shall first discuss the general case before considering two special cases in detail. The first special case has the plasma density constrained by pressure balance, as occurs in astrophysical and recent experimental contexts (Markevitch and Vikhlinin, 2007; Meinecke et al., 2022). The second special case will be the simpler situation in which the density is constant, corresponding to the analysis performed in Bell et al. (2020).
@@ 7.1 General theory
Let us consider monotonic profiles such that $\partial_{z}\mathcal{M}\neq 0$ everywhere, $\mathcal{M}$ having been defined in (17) but with $\epsilon=0$. Then one can formally parameterize the inverse function $z(\mathcal{M})$ so that all functions can be considered functions of $\mathcal{M}$. Suppose further that
@@
where $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are arbitrary values at which boundary conditions can be applied. One notes that (66) manifestly respects the affine symmetry of the original equation. For (66) to be physically relevant, though, it must be the case that $\mathcal{M}(z)\geq 0$; a sufficient condition to ensure positivity is that $g\sim A/\mathcal{M}$ with $A>1$ as $\mathcal{M}\to 0$, as shown in Appendix F. Realistic friction coefficients do indeed have this property, as we shall see in Sec. 7.3 and Sec. 7.4.
@@ 7.2 Magnetization staircases as a general class of solutions
Let us now consider a high-$\beta_{0}$ plasma. Although not obvious from (66), in this limit, $\mathcal{M}(z)$, and thus $T(z)$, naturally forms a staircase structure. To see this more easily, note that $G_{1}\sim O(\delta^{-1})$ and $G_{2}\sim O(1)$ with respect to the small parameter $\delta\sim 1/\beta_{0}$ [see (60), or more simply, see that $F_{T,1}$ is significantly larger than any other coefficient in Fig. 9 when $\beta_{0}$ is large]. Hence, (63) has the general abstract form
@@
where $G$ is a nominally $O(1)$ function. It is well established (Bender and Orszag, 1978) that the solutions to such equations can exhibit boundary layers when $\delta\to 0$; for (67) such boundary layers will occur where $G(y)=0$.
@@
Figure 11: Solutions of (67) when $G(y)$ takes the form shown in each respective inset. All solutions have $|\delta|=0.01$. The ‘arbitrary units’ (a.u.) designation on the $x$-axis emphasizes that, due to affine symmetry, there is formally no scale to the $x$-dependence of the solutions. All solutions satisfy $y(0)=0$, with the other boundary condition $y^{\prime}(0)$, which simply controls the horizontal scale of the solution, adjusted for each case to fit the pertinent behaviour on the same axis. The functional forms for the plots shown in (a) and (b) can be derived analytically, as shown in Appendix G.
The basis for this conjecture is demonstrated in Fig. 11, which shows solutions of (67) for a polynomial $G(y)$. The derivation of (68) and the analytical solution for certain special cases of $G(y)$ are presented in Appendix G. Let us now demonstrate the role that these staircase solutions play in determining the globally stable temperature profiles for isobaric and constant-density plasmas.
@@ 7.3 Isobaric density profile with Chapman–Enskog friction
Using known asymptotics (111) of the Lorentz friction coefficients, it is straightforward to show that $\mathcal{M}g\to 8/5$ as $\mathcal{M}\to 0$. Hence, the temperature profile is guaranteed to be positive everywhere (Appendix F). To obtain a simple analytical approximation for the solution to (66), it is reasonable to take
@@
Note, importantly, that a geometrical-optics description of this parameter regime is not valid because (39) predicts that $k_{z,\text{max}}L\sim\mathcal{M}_{*}\ll 1$ when $\mathcal{M}_{*}^{2}\beta_{0}\sim 1$ and $\mathcal{M}_{*}\ll 1$. The continuation of the root line to small $\beta_{0}$ can be computed using the $\mathcal{M}\to\infty$ limit of the friction coefficients (112) to give
@@
As shown in Appendix G, the solution to (66) can be computed analytically for $g(\mathcal{M})$ given by (74):
@@ 8 Back-reaction on the background temperature profile
The appropriate lowest-order expressions for $\eta$, $v_{N}$, and $\mathcal{D}_{A}$ are given in (27) and (35b). It is important to note that, when combined with (36), (86b) manifestly conserves the total energy of the electron MHD system of equations (the appropriate expressions for energy conservation are presented in Appendix H).
@@ 8.1 Frictional cooling
$\frac{\mathcal{M}\beta_{0}}{2\alpha_{\perp}}\left(\sqrt{\beta_{\wedge}^{2}+\Delta_{\beta}^{2}}+\beta_{\wedge}\right)<L_{T}L_{I}^{-1}<\frac{\mathcal{M}\beta_{0}}{2\alpha_{\perp}}\left(\sqrt{\beta_{\wedge}^{2}+\Delta_{\beta}^{2}}-\beta_{\wedge}\right).$ (92)
@@ 8.2 Reduced heat flux
$\frac{q_{z}}{q_{0}}=\kappa_{\parallel}+\frac{\epsilon^{2}}{2}\left(\frac{\Delta_{\beta}^{2}}{2\alpha_{\perp}}+\frac{\beta_{\wedge}}{\mathcal{M}\beta_{0}}\frac{L_{T}}{L_{I}}-\Delta_{\kappa}\right)\mathcal{I},\quad q_{0}=\frac{nTv_{t}^{2}\tau_{ei}}{L_{T}}.$ (93)
Hence, we see that the net effect of the instability on the heat flux results from the competition of three terms. The first two $O(\epsilon^{2})$ terms are associated with the Ettingshausen effect [^5] For detailed discussions of the Ettingshausen effect, see, e.g., Chittenden and Haines (1993) and Kolmes et al. (2021a)., which is the additional heat flux [beyond the standard enthalpy flux (Epperlein and Haines, 1986)] carried by faster moving, less collisional electrons whose directional symmetry is broken with a mean flow. The first of these terms is always positive and therefore always enhances the heat flux; in contrast, the second term can reduce the heat flux when $L_{I}$ and $L_{T}$ are oppositely oriented and can even overcome the first term if $L_{T}L_{I}^{-1}$ is sufficiently negative (meaning that $\mathcal{I}$ is sufficiently sharply peaked):
@@
where the final approximation is for $\mathcal{M}\ll 1$. This heat-flux-reduction mechanism can be readily achieved in the high-$\beta_{0}$, low-$\mathcal{M}$ regime in which the temperature staircases discussed in Sec. 7 also form, since in this limit the right-hand side of (94) goes to zero as $25/\beta_{0}$ [see (75)]. The third $O(\epsilon^{2})$ term in (93), which is always negative, is the reduction of the effective conductivity due to the transverse magnetic field perturbations generated by the instability causing the temperature gradient and the total magnetic field to become misaligned.
@@ 8.3 Marginally stable heat flux
Finally, let us suppose that the temperature profile is in a globally marginally stable state such that $\left\langle\mathcal{D}_{A}\right\rangle=0$ (Sec. 7). Dynamically, one expects an arbitrary temperature profile to be driven towards such a state on the instability timescale, which can be faster than the conduction timescale (see Fig. 10). This is because, as shown in Fig. 9, the instability growth rate $\mathcal{D}_{A}$ is an increasing function of temperature (and conversely, the damping rate is a decreasing function of temperature). Energy conservation (86b) then implies that higher temperatures are increasingly cooled by the instability (thereby decreasing $\mathcal{D}_{A}$) while lower temperatures are increasingly heated (thereby decreasing $\mathcal{D}_{A}$) to create temperature plateaus separated by transition regions where the temperature profile has remained unchanged because $\mathcal{D}_{A}\approx 0$ initially. The result is a profile that has $\mathcal{D}_{A}=0$ everywhere.
@@
$\frac{\mathcal{Q}_{z}}{\kappa_{\parallel}q_{0}}=1-\epsilon^{2}\Gamma\mathcal{I},$ (95)
@@
Here we have imposed pressure balance for simplicity; the general expression is obtained by replacing $5/2$ with $L_{T}\partial_{z}\mathcal{M}/\mathcal{M}$ in the second term. As in (93), we see that the intensity gradient is capable of reducing the total heat flux. In this case, the two mechanisms available are the Ettingshausen mechanism discussed in Sec. 8.2 and the frictional cooling discussed in Sec. 8.1 (now coupled to the heat flux by imposing global marginal stability). Both are controlled by the lengthscale ratio $L_{T}L_{I}^{-1}$. As shown in Fig. 15, making $L_{T}L_{I}^{-1}$ negative causes the heat flux to be reduced ($|\mathcal{Q}_{z}|<\kappa_{\parallel}q_{0}$) over a large region of parameter space [^6] One can have reduced heat flux even with $L_{T}L_{I}^{-1}>0$ provided that $\beta_{0}$ and $\mathcal{M}$ are both sufficiently large ($\beta_{0}\mathcal{M}\gtrsim L_{T}L_{I}^{-1}$) such that $\Delta_{\kappa}$ dominates (96)..
@@
Indeed, the flux-reduction factor $\Gamma$ can be made arbitrarily large by making $L_{T}L_{I}^{-1}$ increasingly negative. To make this more quantitative, let us consider the weakly magnetized (small-$\mathcal{M}$) limit and approximate all transport coefficients by their lowest-order asymptotic limits (111). Then one has
@@
or, equivalently, using (97)
$\frac{L_{T}}{L_{I}}\geq\frac{\beta_{0}}{0.423}\left(\frac{\kappa_{\parallel}+\kappa_{\perp}}{2\kappa_{\parallel}}\right)\gtrsim 2\beta_{0},$ (100)
@@
Thus, the collisional whistler instability is capable of reducing the electron heat flux in principle, somewhat similar to the collisionless whistler instability (Levinson and Eichler, 1992; Pistinner and Eichler, 1998; Komarov et al., 2018; Roberg-Clark et al., 2018). The persistence time $\tau$ of the temperature profile is then nominally lengthened by the same factor $\tau\sim\tau_{\kappa}/|1-\epsilon^{2}\Gamma\mathcal{I}|$ until it is ultimately set by the persistence time of the intensity profile itself, which in turn is set by advection and refraction (i.e., evolving the $z$ and the $k_{z}$ dependence of $\mathcal{W}$, respectively). The advection-limited persistence time is expected to be comparable to $\tau$ (and thus not a limiting factor) since the total Poynting flux that accounts for the intensity advection [the term in square brackets in (157)] and the modified heat flux $\mathcal{Q}_{z}$ used to estimate $\tau$ only differ by subdominant terms. The refraction timescale, however, is difficult to estimate given that it inherently involves breaking the quasi-eikonal ansatz (88), requiring one to reconsider the full phase-space dynamics of $\mathcal{W}$. This is beyond the scope of the present work.
Let us conclude this section with a brief word of caution regarding the bound obtained in (100). The inequality (100) represents a stricter requirement on $L_{I}$ compared to (94) because (100) requires second-order effects to become comparable to the lowest-order heat flux, whereas (94) results from comparing two second-order terms. Clearly, such an occurrence would also indicate that the perturbative approach underlying (87) has potentially become invalid; if $L_{T}L_{I}^{-1}$ is too large, even an infinitesimal intensity will grow quickly in space and become relatively large elsewhere. Therefore, our analysis here is merely a simple first attempt to understand what is required of $\mathcal{I}(z)$ to greatly reduce the heat flux assuming quasilinear theory holds for all $L_{T}L_{I}^{-1}$. Future investigations can be conducted to see how nonlinear physics modifies this constraint. Conversely, it is likely that the heat-flux suppression by the collisional whistler instability will be a small effect for high-beta plasmas within the strict validity of our fluid slab model.